OpenStudy (anonymous):

Solve the following equation. Please show work. -8/9x=48

jimthompson5910 (jim_thompson5910):

Multiply both sides by -9/8 to get -8/9x = 48 (-9/8)(-8/9x) = 48(-9/8) x = -54